OpenStudy (anonymous):

The mean of 8 numbers is 54.75. What would the new mean be if the numbers 94 and 102 are added to the data set?

OpenStudy (anonymous):

I thought that you would take that mean and add 94 and 102 to it and then divide that by 10 and I ended up getting 25.075 but that isn't one of my options.

OpenStudy (cwrw238):

total after adding 94 and 102 = 8 * 54.75 + 94 + 102 new mean = this total / 10

OpenStudy (anonymous):

54.75+102+94=250.75 250.75/10=25.075

OpenStudy (cwrw238):

you should multiply 54.75 by 8 to get the 'old' total

OpenStudy (anonymous):

Ohh i didn't multiply

OpenStudy (anonymous):

Okay I get it now thanks.

OpenStudy (cwrw238):

yw